rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6 0 g 0 G rgb 0.2 0.2 0.6

0 g 0 G

0 g 0 G0 g 0 G rgb 0.2 0.2 0.6Planeettojen suunnittelu0 g 0 G

Sampo Smolander

Esa Peki

rgb 0.2 0.2 0.6 10 g 0 G rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6 0 g 0 G

rgb 0.2 0.2 0.6 Why do this?

0 g 0 G

0 g 0 G0 g 0 G Photons, coming from sky and hitting the forest 0 g 0 G0 g 0 G can have 3 different fates:

0 g 0 G0 g 0 G 1: Absorbed in soil 0 g 0 G0 g 0 G 2: Absorbed in forest canopy

0 g 0 G0 g 0 G ⟶ provide energy for photosynthesis

0 g 0 G0 g 0 G 3: Reflected back to sky

0 g 0 G0 g 0 G ⟶ these carry information about the forest structure

rgb 0.2 0.2 0.6Sampo Smolander 2 0 g 0 G

rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6rgb 0.2 0.2 0.6 0 g 0 G

0 g 0 G0 g 0 G especially, if happen to be seen by a satellite

rgb 0.2 0.2 0.6Sampo Smolander 3 0 g 0 G